# QCircuitNet: A Large-Scale Hierarchical Dataset for Quantum Algorithm Design

### v1

@@ 5.2 Fine-tuning on QCircuitNet
We compare the performance of Llama3-8B before and after fine-tuning with case studies. Take oracle construction of Bernstein-Vazirani Problem as an example, we observed that before fine-tuning, the model would indiscriminately apply CX gates to all qubits. After fine-tuning, it began to selectively apply CX gates to qubits with ’1’s in the secret string. In some cases, the positions were still counted incorrectly; however, in certain instances, the model accurately identified all the positions for applying the CX gates, which is highly impressive. This improvement significantly contributed to higher scores, suggesting that the model is starting to learn the pattern for constructing certain oracles through fine-tuning. Regarding the interesting performance decrease on Clifford and universal random circuits, we conducted additional experiments on temperature and refer to Appendix C.2 for more details.
@@ 6 Conclusions and Future Work
In this paper, we propose QCircuitNet, the first comprehensive, structured universal quantum algorithm dataset and quantum circuit generation benchmark for AI models. This framework formulates quantum algorithm design from the programming language perspective and includes detailed descriptions and implementation of most established and important quantum algorithms / primitives, allowing for automatic verification methodologies. Benchmarking of QCircuitNet on up-to-date LLMs is systematically conducted. Fine-tuning results also showcase the potential of QCircuitNet as a training dataset. As shown by these benchmarking and fine-tuning results, QCircuitNet helps guide LLMs for reasoning and implementation. As we gradually extend the dataset, it will hopefully contribute to interactive quantum algorithm design.
We also highlight the challenges and opportunities as follows, ranging from scalability on the quantum algorithm side to data contamination on the AI side:
Scalability of the Approach.
Our framework is designed to scale with increasing qubit numbers and support complex quantum algorithms as long as they are efficiently implementable with polynomial gates. The dataset generation scripts are written in a generic way which can be easily extended to arbitrarily large qubit numbers. The implementation of Generalized Simon’s Problem mentioned in Section 4.1.2 showcases the compatibility of our framework with more complex algorithms. The main bottleneck for scalability in the pipeline lies in the simulation process in verification function. For simplicity of demonstration and consideration of hardware noise, all the verification functions were run with classical simulations in our experiments. But the APIs we implemented are compatible with IBM hardware and can be easily adapted to quantum computers, which reduces the cost of classical simulation. As quantum hardware becomes widespread and the capabilities of large language models continue to advance, the dataset can be easily expanded to accommodate scaling.
Data Contamination in AI Learning.
We observe a performance separation between writing general Qiskit codes and explicit gate-level circuits in QASM. Since Qiskit provides detailed tutorial with general codes for several algorithms, this may imply a data contamination phenomenon where LLMs rely on memorization and retrieval rather than genuine algorithm design (see Appendix C.4 for more details). Similarly, current benchmarks for AI code generation and syntax learning may also suffer from this unseen bias. Our dataset, based on QASM files created from scratch, may help circumvent this issue and serve as a stable and fair method for benchmarking AI syntax learning.
@@
  • QCircuitNet is a benchmarking dataset for LLMs. It is of general interest to extend benchmarking to training, which will help LLMs better maneuver quantum algorithm design. We have implemented advanced algorithms such as the Generalized Simon’s Problem, but this in general needs implementations of more advanced algorithms to make it a more meaningful training dataset.
  • Since quantum algorithms have fundamental difference from classical algorithms, novel fine-tuning methods to attempt quantum algorithm design and quantum circuit implementation, or even developments of new quantum algorithms by LLMs are solicited.
  • Currently, variational quantum algorithms (Cerezo et al., 2021) can already be implemented on near-term NISQ machines (Preskill, 2018). It would be also of general interest to extend QCircuitNet to contain the design and implementation of variational quantum algorithms.
@@ Acknowledgements
RY and TL were supported by the National Natural Science Foundation of China (Grant Numbers 92365117 and 62372006), and The Fundamental Research Funds for the Central Universities, Peking University. We would like to thank Hengli Li and Zihao Wang for their valuable discussions.
Appendix A Details of QCircuitNet
The QCircuitNet Dataset is available on GitHub at the following link: https://github.com/EstelYang/QCircuitNet_Dataset
[…]
We expect to extend QCircuitNet under this general structure.
@@ A.1 Format
In this subsection, we provide concrete examples to illustrate the different components of QCircuitNet. We use the case of Simon’s Problem throughout the demonstration to achieve better consistency. For further details, please check the code repository.
@@
    def check_model(qasm_string, code_string, n):
    """Check the Simon model."""
@@
    t = 1
    with open(f"test_oracle/n{n}/trial{t}/oracle.inc", "r") as file:
    oracle_def = file.read()
@@
    return -1
@@
    exec(code_string, globals())
    aer_sim = AerSimulator()
    total_success = 0
    total_fail = 0
    t_range = min(10, 4 ** (n - 2))
    shots = 10
    for t in range(1, 1 + t_range):
    print(f" Running Test Case {t}")
    with open(f"test_oracle/n{n}/trial{t}/oracle.inc", "r") as file:
    oracle_def = file.read()
    full_qasm = plug_in_oracle(qasm_string, oracle_def)
    circuit = loads(full_qasm)
    with open(f"test_oracle/n{n}/trial{t}/oracle_info.txt", "r") as file:
    content = file.read()
    match = re.search(r"Secret string: ([01]+)", content)
    if match:
    secret_string = match.group(1)
    else:
    raise ValueError("Secret string not found in the file.")
    cnt_success = 0
    cnt_fail = 0
    for shot in range(shots):
    prediction = run_and_analyze(circuit.copy(), aer_sim)
    if not isinstance(prediction, str):
    raise TypeError("Predicted secret string should be a string.")
    if prediction == secret_string:
    cnt_success += 1
    else:
    cnt_fail += 1
    print(f" Success: {cnt_success}/{shots}, Fail: {cnt_fail}/{shots}")
    total_success += cnt_success
    total_fail += cnt_fail
    print(f"Total Success: {total_success}; Total Fail: {total_fail}")
    return total_success / (total_fail + total_success)
@@
    print(f"Error: {e}")
    return -1
@@
A.2 Discussion of more Tasks
@@
In the "Quantum Information" section of the "Algorithm Design" task, we have also implemented three important quantum information protocols: Quantum Teleportation, Superdense Coding, and Quantum Key Distribution (BB84). A brief introduction to these protocols can be found in Appendix B. We did not include the experiments for these protocols as they involve communication between two parties, which is challenging to characterize with a single OpenQASM 3.0 file. We recommend revising the post-processing function as a general classical function to schedule the communication and processing between different parties specifically for these protocols. The fundamental quantum circuits and processing codes are provided in the repository.
A.3 Datasheet
Here we present a datasheet for the documentation of QCircuitNet.
Motivation
  • For what purpose was the dataset created? It was created as a benchmark for the capability of designing and implementing quantum algorithms for LLMs.
  • Who created the dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)? The authors of this paper.
  • Who funded the creation of the dataset? We will reveal the funding resources in the Acknowledgement section of the final version.
Composition
  • What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)? The dataset comprises problem description, generation code, algorithm circuit, post-processing function, oracle / gate definition, verification function, and dataset creation script for various quantum algorithms.
  • How many instances are there in total (of each type, if appropriate)? The dataset has 5 algorithms for oracle construction task and 10 algorithms for algorithm design task used for experiments. There are 3 quantum information protocols and additional problem encoding tasks not included for experiments.
  • Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set? The dataset contains instances with restricted qubit numbers due to the current scale of real quantum hardware.
  • What data does each instance consist of? Qiskit codes, OpenQASM 3.0 codes, python scripts, and necessary text information.
  • Are relationships between individual instances made explicit? Yes, the way to create different instances are clearly described in Appendix A.1.
  • Are there recommended data splits? Yes, we recommend splitting the data according to different algorithms in algorithm design task.
  • Are there any errors, sources of noise, or redundancies in the dataset? There might be some small issues due to the dumping process of Qiskit and programming mistakes (if any).
  • Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)? The dataset is self-contained.
  • Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor-patient confidentiality, data that includes the content of individuals’ non-public communications)? No.
  • Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? No.
Collection Process
  • How was the data associated with each instance acquired? The data is created by first composing Qiskit codes for each algorithm and then converting to OpenQASM 3.0 files using Qiskit.qasm3.dump function, with additional processing procedure.
  • What mechanisms or procedures were used to collect the data (e.g., hardware apparatuses or sensors, manual human curation, software programs, software APIs)? Manual human programming and Qiskit APIs.
  • Who was involved in the data collection process (e.g., students, crowd workers, contractors), and how were they compensated (e.g., how much were crowd workers paid)? Nobody other than the authors of the paper.
  • Over what timeframe was the data collected? The submitted version of the dataset was created in May and June 2024.
Uses
  • Has the dataset been used for any tasks already? It has been used in this paper to benchmark LLM’s ability for quantum algorithm design.
  • Is there a repository that links to any or all papers or systems that use the dataset? The only paper which uses the dataset for now is this paper.
Distribution
  • Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created? Yes, the dataset will be made publicly available on the Internet after the review process.
  • How will the dataset be distributed (e.g., tarball on website, API, GitHub)? It will be distributed on the GitHub platform.
  • Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)? The dataset is distributed under CC BY 4.0.
  • Have any third parties imposed IP-based or other restrictions on the data associated with the instances? No.
  • Do any export controls or other regulatory restrictions apply to the dataset or to individual instances? No.
Maintenance
  • Who will be supporting/hosting/maintaining the dataset? The authors of this paper.
  • How can the owner/curator/manager of the dataset be contacted (e.g., email address)? The email for contact will be provided after the review process.
  • Is there an erratum? Not at this time.
  • Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)? Yes, it will be continually updated.
  • If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? Yes, they can do so with the GitHub platform.
A.4 Copyright and Licensing Terms
This work is distributed under a CC BY 4.0 license. The implementation of the code references open-source projects such as Qiskit, QuantumKatas, Cirq, and NWQBench. We bear responsibility in case of violation of rights.
Appendix B Additional Preliminaries for Quantum Computing and Quantum Information

### v2

@@ 4.2 Fine-tuning on QCircuitBench
We observe that the Llama3 model demonstrates the most notable improvement on Grover’s algorithm after fine-tuning, with the verification score increased by 0.1900. Case studies on the Bernstein-Vazirani oracle reveal that, before fine-tuning, the model would indiscriminately apply CX gates to all qubits. After fine-tuning, it begins to selectively apply CX gates to qubits corresponding to ‘1’s in the secret string. While some counting errors persist, the model occasionally identifies all correct positions, demonstrating a marked improvement. This suggests that fine-tuning enables the model to internalize structural patterns in oracle construction, leading to improved performance across tasks.
Regarding the interesting performance decrease on Clifford and universal random circuits, we conducted additional experiments on temperature and entropy. We refer the readers to Appendix LABEL:appendix-finetune-random for more details. We also performed fine-tuning on the quantum algorithm design task, which is included in Appendix LABEL:appendix-finetune-algo.
@@ 5 Conclusions and Future Work
In this paper, we propose QCircuitBench, the first comprehensive, structured quantum algorithm dataset and quantum code generation benchmark for AI models. We highlight the main takeaways as follows:
  • Novelty and Impact: QCircuitBench introduces the first comprehensive benchmark specifically tailored for LLM-driven quantum algorithm design, addressing a critical research gap at the intersection of quantum computing and AI, with implications for both communities.
  • Dataset Design: The dataset defines a general framework which formulates the key components of quantum algorithm design in a manner suitable for LLMs, with careful considerations for various subtle challenges. Its code generation perspective, modular and extensible structure, automatic verification functions, and broad coverage of diverse quantum algorithms together establish a rigorous benchmark for evaluating the capabilities of LLMs in quantum algorithm design. The framework is scalable to larger qubit numbers and can incorporate more advanced algorithms that are polynomially implementable.
  • Extensive experiments: Our experiments show that QCircuitBench poses significant challenges to state-of-the-art LLMs, making it a valuable benchmark for AI models. We identify three predominant error types: (1) improvisation errors, where models hallucinate or deviate from task constraints such as syntax or required structure; (2) counting errors, especially in qubit or gate indices; and (3) data contamination, likely stemming from prior exposure to known implementations. Despite these challenges, fine-tuning experiments demonstrate early promise in helping models internalize structural patterns. These insights position QCircuitBench not only as a benchmark, but also as a path for future model development.
  • Clarity and Reproducibility: We release the dataset in a fully reproducible format, accompanied by detailed documentation, executable code and transparent benchmarking protocols. This ensures that researchers can effortlessly use, validate, and extend the benchmark, with the aim of contributing to open scientific efforts.
@@
  • While QCircuitBench is designed as a benchmark, its structure also supports training objectives. It is of general interest to extend benchmarking to training, which will help LLMs better maneuver quantum algorithm design. We have implemented advanced algorithms such as the Generalized Simon’s Problem, but further coverage of more advanced algorithms will make it more impactful.
  • As quantum algorithms differ fundamentally from classical ones, novel fine-tuning methods and model architectures may be required. These could focus on quantum-specific inductive biases, hybrid neural-symbolic pipelines, or even autonomous quantum algorithm discovery via LLMs, potentially enabling generative models to contribute novel quantum methods.
@@ Acknowledgements
We would like to thank Zihang Li, Kecheng Liu, Ruihua Liu, Shaozhen Liu, Zhixin Song, Kai Wang, Mengfan Yuan, Lu Zhang, Shuo Zhang, Xingyu Zhao for participating in the human study. Special thanks to Hengli Li and Zihao Wang for their valuable discussions.
This work was supported by the National Natural Science Foundation of China (Grant Number 92365117).
Appendix A Details of QCircuitBench
The QCircuitBench Dataset, along with detailed documentation, benchmarking scripts, and Croissant metadata, is available on GitHub (https://github.com/EstelYang/QCircuitBench) and Harvard Dataverse (https://doi.org/10.7910/DVN/ZC4PNI). We recommend visiting the GitHub repository for the latest updates.
The following structure represents the default Qiskit-based implementation of QCircuitBench. The separate directory named Cirq_Version mirrors this structure to provide equivalent implementations using Cirq.
QCircuitBench
    Algorithm Design ........................................................................................All data and code for the quantum algorithm design task
    Quantum Computing ........................................................................................Universal quantum computing algorithms
    Quantum Information ........................................................................................Quantum information protocols
    Variational Quantum Algorithms ........................................................................................Variational quantum algorithms
    Oracle Construction ........................................................................................All data for the oracle construction task
    Quantum Logic Synthesis ........................................................................................Textbook-level and advanced oracles
    Problem Encoding ........................................................................................Oracles encoding application scenarios
    Random Circuits ........................................................................................All data for the random circuit synthesis task
    Clifford ........................................................................................Random circuits with the Clifford gate set
    Universal ........................................................................................Random circuits with the universal gate set
    Cirq Version ........................................................................................Implementation using Cirq
In each subdirectory, there is a folder for each specific algorithm. For instance, the folder structure for Simon’s algorithm is as follows:
Algorithm Design
    Quantum Computing
    simon ........................................................................................All data for the Simon’s Problem
    simon-dataset.py ........................................................................................Dataset creation script
    simon-generation.py ........................................................................................Qiskit generation code
    simon-post-processing.py ........................................................................................Post-processing function
    simon-utils.py ........................................................................................Utility functions for verification
    simon-verification.py ........................................................................................Verification function
    simon-description.txt ........................................................................................Problem description
    simon-verification.txt ........................................................................................Verification results of the data points
    full circuit ........................................................................................Raw data of quantum circuits
    simon-n2
    simon-n2-s11-k11.qasm ........................................................................................Full circuit for a concrete setting
    simon-n3
    simon-n3-s011-k001.qasm
    simon-n3-s011-k101.qasm
    test oracle ........................................................................................Extracted oracle definitions
    n2
    trial1
    oracle.inc ........................................................................................Oracle definition as a .inc file
    oracle-info.txt ........................................................................................Oracle information (such as key strings)
    n3
    trial1
    oracle.inc
    oracle-info.txt
    trial2
    oracle.inc
    oracle-info.txt
    simon-n2.qasm ........................................................................................Algorithm circuit for model output
    simon-n3.qasm
We expect to extend QCircuitBench under this general structure.
@@ A.1 Format
In this subsection, we provide concrete examples to illustrate the different components of QCircuitBench. We use the case of Simon’s Problem throughout the demonstration to achieve better consistency. For further details, please check the code repository.
@@
    def check_model(qasm_string, code_string, n, t=1):
    """Check the Bernstein-Vazirani model."""
@@
    qasm_syntax = -1
    code_syntax = -1
    result_score = 0.
    gate_count_ratio = float(’nan’)
    shot_ratio = float(’nan’)
    time_ratio = float(’nan’)
    # QASM syntax verification
@@
    print("QASM syntax error detected, using ground truth.")
    qasm_syntax = 0
    dire_gt = "lm_eval/tasks/QuantumAlgorithm/Circuit/simon"
    with open(f"{dire_gt}/simon_n{n}.qasm", "r") as file: # model generated is wrong, use ground truth
    qasm_string = file.read()
    else:
    qasm_syntax = 1
    # Post-Processing code verification
@@
    local_vars = {}
[…]
    exec(code_string, globals(), local_vars)
    run_and_analyze_func = local_vars[’run_and_analyze’]
    code_syntax = 1
    print("Post-processing code loaded successfully.")
@@
[…]
    run_and_analyze_func = ground_truth_run_and_analyze
    code_syntax = 0
    if qasm_syntax == 1 and code_syntax == 1: # Only check the case : model has at least one correct
    gate_count_ratio, shot_ratio, time_ratio = efficiency_check(qasm_string, dire_gt, code_string, run_and_analyze_func, ground_truth_run_and_analyze, dire, n)
    try:
    result_score = execute_test_cases(qasm_string, run_and_analyze_func, n)
    except Exception as e:
    print(f"Post-processing running-time error: {e}")
    code_syntax = 0
    return qasm_syntax, code_syntax, result_score, gate_count_ratio, shot_ratio, time_ratio
@@
A.2 Visualization of reference circuits
In this subsection, we present visualizations of reference quantum circuits for each of the 25 algorithms across all three task suites.
Figure 5: Visualization of Reference Quantum Circuits for Quantum Algorithm Design Task (I).
Figure 6: Visualization of Reference Quantum Circuits for Quantum Algorithm Design Task (II).
Figure 7: Visualization of Reference Quantum Circuits for Quantum Algorithm Design Task (III).
Figure 8: Visualization of Reference Quantum Circuits for Quantum Algorithm Design Task (IV).
Figure 9: Visualization of Reference Quantum Circuits for Quantum Algorithm Design Task (V).
Figure 10: Visualization of Reference Quantum Circuits for Random Circuit Synthesis Task.
Figure 11: Visualization of Reference Quantum Circuits for Oracle Construction Task (I).
Figure 12: Visualization of Reference Quantum Circuits for Oracle Construction Task (II).
A.3 Human study
Here we present the instructions for the human study reported in Section 4.1.
Figure 13: Instructions for human study.
A.4 Discussion of more tasks
@@
In Section 3.1.2, we have also implemented three important quantum information protocols: Quantum Teleportation, Superdense Coding, and Quantum Key Distribution (BB84). A brief introduction to these protocols can be found in Appendix B. We did not include the experiments for these protocols as they involve communication between two parties, which is challenging to characterize with a single OpenQASM 3.0 file. We recommend revising the post-processing function as a general classical function to schedule the communication and processing between different parties specifically for these protocols. The fundamental quantum circuits and processing codes are provided in the repository.
Appendix B Preliminaries for Quantum Computing and Quantum Information
In this section, we will introduce necessary backgrounds for quantum computing related to this paper. A more detailed introduction to quantum computing can be found in the standard textbook by nielsen2000book.
Quantum States.
Here $|\cdot\rangle$ represents a column vector, also known as a ket state. The tensor product of two quantum states $|\phi_{1}\rangle=\sum_{i=0}^{N-1}\alpha_{i}|i\rangle$ and $|\phi_{2}\rangle=\sum_{j=0}^{M-1}\beta_{j}|j\rangle$ with $M=2^{m}$, $m\in\mathbb{N}$ is defined as
Quantum Oracles.
$U_{f}|x\rangle|z\rangle=|x\rangle|z\oplus f(x)\rangle,$ (3)
Quantum Gates.